OpenStudy (anonymous):

help finding the common denominator of each group of fractions 1/2 and 1/3 besides 6 which i got ?

OpenStudy (anonymous):

basically any number whose factor is 2 and 3 would be a common denominator. 12,24,60,... etc.

jimthompson5910 (jim_thompson5910):

The LCD is 6 Any common denominator is then a multiple of this LCD So other common denominators are: 12, 18, 24, 36 Since they are all multiples of the LCD 6

OpenStudy (anonymous):

srry it says find the least common denominator besides 6 ?

OpenStudy (anonymous):

and it should include more than one fraction

jimthompson5910 (jim_thompson5910):

Well the only LCD is 6 and I only see two fractions. Is there something missing?

OpenStudy (anonymous):

wats the fraction ?

OpenStudy (anonymous):

is it 5/6 ?

jimthompson5910 (jim_thompson5910):

If you want to add 1/2 and 1/3 then.... 1/2+1/3 = 3/6+2/6 = (3+2)/6=5/6 So 1/2+1/3=5/6

OpenStudy (anonymous):

thx